OpenStudy (anonymous):

@johnweldon1993 i need help again i tried the substition thing u showed me but i got it wrong

OpenStudy (anonymous):

The Tran family and the Chen family each used their sprinklers last summer. The water output rate for the Tran family's sprinkler was 20L per hour. The water output rate for the Chen family's sprinkler was 30L per hour. The families used their sprinklers for a combined total of 40 hours, resulting in a total water output of 950 . How long was each sprinkler used?

OpenStudy (johnweldon1993):

no problem :) whats the question?

OpenStudy (johnweldon1993):

Okay....Let me see how you set it up...

OpenStudy (anonymous):

i did 20x+30y=950 and x+y=40

OpenStudy (johnweldon1993):

Okay that is correct.....and what did you do from there? *I'm pretty much going to make you list what you did and see what went wrong :)

OpenStudy (anonymous):

y=-x+40 then 20x+30(-x+40)=950

OpenStudy (johnweldon1993):

good so far...

OpenStudy (anonymous):

then -10x=830 then i got -83

OpenStudy (anonymous):

but it looks wrong

OpenStudy (johnweldon1993):

ahh there we go.....what is 30 times 40?

OpenStudy (anonymous):

120 i subtracted it from 950 to get the 830

OpenStudy (johnweldon1993):

30 times 40 is not 120.... 40 times 3...is 120 40 times 30...is...?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

ohhhhh 1200

OpenStudy (johnweldon1993):

there you go :)

OpenStudy (johnweldon1993):

now do it again and what do you get for x?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

so y is 15

OpenStudy (johnweldon1993):

perfect...and now ....yes y would be 15...check them to make sure they're correct in that first equation...

OpenStudy (anonymous):

i did they were

OpenStudy (johnweldon1993):

Then you're all set! ... Good job :)

OpenStudy (anonymous):

what about this one A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large box and a small box is 95 pounds. The truck is transporting 65 large boxes and 55 small boxes. If the truck is carrying a total of 5825 pounds in boxes, how much does each type of box weigh?

OpenStudy (johnweldon1993):

set it up for me :)

OpenStudy (anonymous):

okay 65x+55y =5825 and x+y=95

OpenStudy (johnweldon1993):

see, you got this!

OpenStudy (anonymous):

U are a life saver lol if i need help with my other topics ill ask u first lol

OpenStudy (johnweldon1993):

lol and I'll do my best to help :)